OpenStudy (anonymous):

i know this is simple but im having a brain fart A motorcycle can travel 70 miles per gallon. Approximately how many gallons of fuel will the motorcycle need to travel 60 km? [1 mile = 1.6 km] 0.2 0.5 0.8 1.4

OpenStudy (anonymous):

@amistre64 60km is 43.75 miles i believe

OpenStudy (anonymous):

@ganeshie8 @satellite73 @SolomonZelman

OpenStudy (amistre64):

what is the conversion rate of mailes to km

OpenStudy (anonymous):

1 mile=1.6km

OpenStudy (amistre64):

70 miles = 1 gallon 1 mile = 1/70 gallons = 1.6 km 1.6 n km = n/70 gallons when does 1.6 n = 60 ?

OpenStudy (anonymous):

37.5

OpenStudy (amistre64):

then we have 37.5/70 gallons used in 60km

OpenStudy (anonymous):

.5

OpenStudy (amistre64):

seems fair to me